this problem I am working on seems so long in solving I was wondering if I need to show each step in this complicated fraction.
Problem is: 6x-1/5 - 2x/3 = 3
what i have so far is this; 3(6x-1)-5*2x/5*3=3
3(6x-1)-10x/5*3=3
(3*6x-3)-10x/5*3=3
(18x-3)-10x/5*3=3
I am not sure if I am on the right track with this one,I would appreciate your help please,
Thank You, Barb Neely

Answer by malakumar_kos@yahoo.com(315)


6x-1/5 - 2x/3 = 3

l.c.m of 3&5 = 15
multiply each fraction by 15 on both sides
we get 3(6x-1)-5(2x) = 3(15)
18x-3-10x = 45
8x = 45+3 = 48
8x = 48 or x = 48/8 = 6 hence x = 6